# Efficient Training-Free Online Routing for High-Volume Multi-LLM Serving

### v2

@@ 2.1 Efficient Performance and Cost Estimation
One significant advantage of using ANNS compared with other training-based approaches is the ease of updating the historical dataset $D$ to adopt various deployment configurations without requiring model retraining. This is particularly advantageous for LLM-serving providers, as it incurs almost no additional deployment overhead. Online systems can naturally collect and record diverse user queries, which can be directly used to maintain and expand the dataset $D$. Although collecting ground-truth performance scores and costs typically requires human or automated evaluation, many providers already integrate such mechanisms (e.g., quality feedback from users) in their systems (chiang2024chatbot, ). Furthermore, widely used ANNS algorithms in practice, such as HNSW [^2] We note that many other choices are interchangeable here, e.g. see https://ann-benchmarks.com., use a graph-based indexing structure, making it substantially more efficient in search complexity (typical $O(\log|D|)$ in practice) compared to traditional instance-based methods like exact K-Nearest Neighbor (KNN, O(|D|)O(|D|)(hu2024routerbench, ), achieving significant speedup in search time.
@@ 4 Evaluation
Benchmarks. We use 3 different benchmarks in our experiments: RouterBench (hu2024routerbench, ), SPROUT (somerstep2025carrot, ), and Open LLM Leaderboard v2 (fourrier2024open, ). RouterBench contains 11 different LLMs, and we randomly sample 10000 queries as the test query dataset and use the remaining data as the historical dataset. Open LLM Leaderboard v2 contains 18 different LLMs, where we similarly sample 10000 queries for the test dataset and use the rest as the historical dataset. SPROUT contains 13 LLMs, and we use the training set as the historical dataset, while the validation and test sets are combined to form the test queries. In the main setting, queries are embedded using bge-base-en-v1.5 bge_embedding . To ensure diversity, we also evaluate with SFR-Embedding-2_R SFR-embedding-2 and gte-Qwen2-1.5B-instruct li2023towards .
Baselines. We compare 8 different routing algorithms, classified into two categories: model-based methods and training-free methods. For model-based methods, following (somerstep2025carrot, ), we train two separate Roberta-based models DBLP:journals/corr/abs-1907-11692 to predict generation performance score and cost, yielding: (i) Roberta-perf-routing, routes each query to the model with the highest predicted performance; (ii) Roberta-cost-routing, routes each query to the model with the greatest available budget. Training-free methods include: (iv) Random routing; (v) Greedy-perf-routing, uses ANNS to select the model with the highest predicted performance; (vi) Greedy-cost-routing, uses ANNS to select the model with the most available budget; (vii) KNN-perf-routing (hu2024routerbench, ), uses KNN to select the model with the highest performance; (viii) KNN-cost-routing (hu2024routerbench, ), uses KNN to select the model with the most available budget; (ix) BatchSplit routing, groups queries into small batches and solves the LP per batch to determine routing. We adopt HNSW (malkov2018efficient, ) as the main ANNS algorithm and set the number of candidate neighbors ($|R_{j}|$) to 5 for both ANNS and KNN. For BatchSplit, we use a mini-batch size of 256 to balance LP computation cost with the low-latency requirements of online routing.
@@ 4.1 Robustness
Routing Overhead & Quality of Historical Data. We evaluate the routing latency of Algorithm˜1 under varying query volumes. As shown in Appendix Table 7, our method incurs negligible decision overhead compared to existing training-free baselines. To assess robustness to data quality, we further consider a noisy setting with label perturbations and an out-of-distribution setting with distribution shift. Results in Appendix Table 8 indicate that our algorithm consistently outperforms all baselines in both settings. See Section˜C.5 for detailed results.
Historical Data & Search Candidates. Further results in Appendix C.6 show that our algorithm remains robust under varying historical data sizes and different numbers of ANNS/KNN candidates.
@@ Appendix A More Experiment Details
MMLU [21] 14042
Hellaswag [63] 10042
GSM8K [9] 7450
ARC Challenge [8] 1470
Winogrande [46] 1267
MBPP [3] 427
MT-Bench [65] 80
@@
MATH Lvl 1-5 [22] 9884
MMLU-PRO [57] 11786
GPQA [45] 541
MUSR [51] 748
RAGBench [18] 1827
Openhermes 2.5 [54] 19455
@@
MMLU-PRO [57] 12032
MUSR [51] 756
MATH Lvl 1-5 [22] 1324
GPQA [45] 1192
BBH [53] 5761
@@
Benchmarks. We evaluate our method on three benchmarks: RouterBench (zero-shot version) [23], SPROUT [50], and Open LLM Leaderboard v2 [16]. All three benchmarks are constructed from multiple dataset sources. Table 2 summarizes the query types in each benchmark, while Table 3 details the LLMs used and their corresponding token costs.
For RouterBench, we identify 13 different data sources in a total of 36497 samples – 6 more than those reported in the original paper – spanning a diverse range of query domains. This benchmark includes 11 different LLMs. Since it does not provide a predefined train/test split, we randomly sample 10000 queries as the test query set and treat the remaining queries as historical data. Note that we do not report per-token costs for models in RouterBench, as [23] does not provide this information. Instead, the benchmark includes precomputed costs for each query across all 11 LLMs, which we directly use in our experiments.
@@
For Open LLM Leaderboard v2, we follow the same setup and data processing as [50], resulting in 21065 samples. The processed benchmark includes 5 different datasets, with each evaluating different aspects of LLM capabilities. This benchmark uses 18 different LLMs, and we randomly sample 10000 queries as the test set and use the remainder as the historical data. Since most evaluations in Open LLM Leaderboard v2 are likelihood-based, the cost is essentially determined by the length of the input. Following [50], we report the cost per input token in Table 3 and compute the total cost of each query based on input token counts.
In the main setting, the queries are embedded using bge-base-en-v1.5 [59]. For the diversity consideration, we additionally evaluate with two different embedding models: SFR-Embedding-2_R [40] and gte-Qwen2-1.5B-instruct [35].
@@
For Model-based methods, we follow the training approach in [50] and train two separate Roberta-based models [36] to predict the performance and cost of each query. The performance prediction model is classification-based, aiming to select the optimal LLM with the highest expected performance, while the cost prediction model is regression-based and estimates the expected cost of an incoming query. This results in 2 different baselines:
@@
  • KNN-perf-routing [23], uses KNN to select the model with the highest performance;
  • KNN-cost-routing [23], uses KNN to select the model with the most available budget;
@@
We adopt HNSW [38] as the main ANNS algorithm and set the number of candidate neighbors ($|R_{j}|$) to 5 for both ANNS and KNN. To assess the robustness of our algorithm under varying candidate set sizes, we further vary this number to 3, 7, and 10. However, many alternative methods, such as DiskANN [25], are interchangeable here, as listed in https://ann-benchmarks.com. For BatchSplit, we use a mini-batch size of 256 to balance LP computation cost with the low-latency requirements of the practical online routing.
@@
Optimization Implementation. We implement all optimization codes using the CVXPY [12] package. For the one-time optimization step in our algorithm, we use the L-BFGS-B solver [66]. For BatchSplit, which involves solving a linear program for each batch, we adopt the HIGHS solver [24].
@@ Appendix B Theoretical Proofs
Definition 1 ($\epsilon$-Net [11]).
@@ B.1 Discussion on Our Theoretical Assumptions
Random arrival order. We assume that the queries arrive in a random order, i.e., they can be picked adversarially, but their order is randomly permuted. Note that this is a weaker requirement than assuming the queries are sampled i.i.d. from an unknown distribution. This is because if the queries are sampled i.i.d., then they also satisfy the random order requirement, and thus our assumption is weaker. Furthermore, the random order model is also a popular assumption in many other online algorithms [39, 11], since it typically allows one to go beyond worst-case hardness and obtain meaningful theoretical guarantees. In practice, we perform robustness studies by changing the query order and observe that our algorithm still outperforms baselines; see Section˜4.1.
@@
Lastly, we remark that similar reasonable lower-bound assumptions have been made in prior work on the online matching problem [19, 11, 39], where it is commonly assumed that bids for the items are small compared to the total budget.
“Estimating query features via similarity” is natural. The core intuition behind most previous literature on predictive LLM routing [23, 43, 56, 50, 52], which trains a model-based predictor on auxiliary datasets to estimate performance and cost of future queries, aligns with Assumption 1. Specifically, these approaches assume that similar queries share similar routing-relevant features. If no such relationship exists, then historical data would be uninformative, and it would be impossible to leverage historical data for effective future routing. In this case, no meaningful inference can be made without directly querying the LLMs, and routing decisions degenerate to random selection.
One may argue that useful routing information can still be obtained by querying only a subset of LLMs. This idea is echoed in works that adopt cascading strategies, where LLMs are queried sequentially based on their observed capabilities on historical data [2, 5]. However, these approaches also fundamentally rely on a related assumption: LLMs that perform well on historical data or benchmarks are likely to perform well on future queries. Furthermore, these methods introduce significant latency and computational overhead, as they may consume substantial token budgets before reaching the optimal LLM. This makes them impractical for high-volume online routing scenarios with tight token budgets and low-latency requirements.
@@ B.3 Proof of Lemma 3
Inspired by the techniques used in [11, 39, 19], we first prove that, under the given setting, the partial performance score can be accurately estimated with high probability. Specifically, for each $i\in[M]$ and $\gamma\in\Phi$, we evaluate the following probability
@@ C.4 Total Budget
C.5 Routing Overhead & Quality of Historical Data
@@
C.6 Historical Data Size & ANNS/KNN Candidates
Historical Data Size. To evaluate the robustness of our algorithm with respect to the size of historical data, we vary the number of historical data points used in ANNS and KNN from 5000 to 25000. For RouterBench, whose maximum available historical data is 26497, we randomly sample the subset of data points from 5000 to 25000. For RouterBench, which contains up to 26497 historical records, we randomly sample subsets within this range. For SPROUT, with a maximum of 30968 records, we similarly sample subsets from 5000 to 25000. For Open LLM Leaderboard v2, which has 11065 historical records, we sample subsets ranging from 5000 to 10000. Note that in this experiment, training-based methods are trained on the full historical dataset to provide a strong baseline for comparison. As shown in Figure 11, the results clearly show that varying the amount of historical data used in ANNS and KNN has minimal impact on performance, with our algorithm consistently and significantly outperforming all baselines across all settings. This demonstrates the robustness and efficiency of our algorithm: even with as few as 5000 historical records, it remains effective for routing.
@@ C.7 Detailed Ablation Study
Impact of embedding models. The choice of embedding model is also important for estimating the performance score and cost of incoming queries. To evaluate the adaptivity of our algorithm to different embedding models used in ANNS, we conduct experiments with two larger embedding models: SFR-Embedding-2_R [40] and gte-Qwen2-1.5B-instruct [35]. As shown in Figure 13, our algorithm consistently outperforms all baselines across all benchmarks, demonstrating its robustness to the choice of embedding model.
@@ C.8 Routing Analysis by Query Types
To gain deeper insights into how our algorithm routes different types of queries, we analyze its routing behavior across the three benchmarks. We categorize query types based on their data sources, as each source is designed to evaluate different LLM capabilities. As shown in Figure 15, 16, and 17, our algorithm consistently achieves superior or competitive routing performance in most query types across benchmarks. These results demonstrate the generalizability and effectiveness of our algorithm in handling diverse query types, which indicates that its performance is not dependent on a narrow subset of tasks.
@@ Appendix D Query-Traffic Considerations in LLM Routing
In the practical LLM-serving systems, an additional important routing factor is query traffic patterns. However, the primary goal of this work is to develop a cost-efficient online routing algorithm, a problem that has gained increasing attention in recent literature [43, 23, 50], especially given the high computational cost of LLMs. To isolate this core algorithmic problem, we adopt a standard assumption widely used in the online algorithms literature [11, 39, 19]: incoming queries arrive in a randomly permuted order. This assumption, discussed in detail in Appendix B.1, allows us to abstract away complex traffic dynamics and focus on the fundamental cost-efficiency routing problem.

### v3

@@ 2.1 Efficient Performance and Cost Estimation
One significant advantage of using ANNS compared with other training-based approaches is the ease of updating the historical dataset $D$ to adopt various deployment configurations without requiring model retraining. This is particularly advantageous for LLM-serving providers, as it incurs almost no additional deployment overhead. Online systems can naturally collect and record diverse user queries, which can be directly used to maintain and expand the dataset $D$. Although collecting ground-truth performance scores and costs typically requires human or automated evaluation, many providers already integrate such mechanisms (e.g., quality feedback from users) in their systems (chiang2024chatbot). Furthermore, widely used ANNS algorithms in practice, such as HNSW [^2] We note that many other choices are interchangeable here, e.g. see https://ann-benchmarks.com., use a graph-based indexing structure, making it substantially more efficient in search complexity (typical $O(\log|D|)$ in practice) compared to traditional instance-based methods like exact K-Nearest Neighbor (KNN, O(|D|)O(|D|)(hu2024routerbench), achieving significant speedup in search time.
@@ 4 Evaluation
Benchmarks. We use 3 different benchmarks in our experiments: RouterBench (hu2024routerbench), SPROUT (somerstep2025carrot), and Open LLM Leaderboard v2 (fourrier2024open). RouterBench contains 11 different LLMs, and we randomly sample 10000 queries as the test query dataset and use the remaining data as the historical dataset. Open LLM Leaderboard v2 contains 18 different LLMs, where we similarly sample 10000 queries for the test dataset and use the rest as the historical dataset. SPROUT contains 13 LLMs, and we use the training set as the historical dataset, while the validation and test sets are combined to form the test queries. In the main setting, queries are embedded using bge-base-en-v1.5 bge_embedding. To ensure diversity, we also evaluate with SFR-Embedding-2_R SFR-embedding-2 and gte-Qwen2-1.5B-instruct li2023towards.
Baselines. We compare 8 different routing algorithms, classified into two categories: model-based methods and training-free methods. For model-based methods, following (somerstep2025carrot), we train two separate Roberta-based models DBLP:journals/corr/abs-1907-11692 to predict generation performance score and cost, yielding: (i) Roberta-perf-routing, routes each query to the model with the highest predicted performance; (ii) Roberta-cost-routing, routes each query to the model with the greatest available budget. Training-free methods include: (iv) Random routing; (v) Greedy-perf-routing, uses ANNS to select the model with the highest predicted performance; (vi) Greedy-cost-routing, uses ANNS to select the model with the most available budget; (vii) KNN-perf-routing (hu2024routerbench), uses KNN to select the model with the highest performance; (viii) KNN-cost-routing (hu2024routerbench), uses KNN to select the model with the most available budget; (ix) BatchSplit routing, groups queries into small batches and solves the LP per batch to determine routing. We adopt HNSW (malkov2018efficient) as the main ANNS algorithm and set the number of candidate neighbors ($|R_{j}|$) to 5 for both ANNS and KNN. For BatchSplit, we use a mini-batch size of 256 to balance LP computation cost with the low-latency requirements of online routing.
@@ 4.1 Robustness
Figure 7: Results on RouterBench when varying the number of historical data points.
Figure 8: Results on RouterBench when varying the number of search candidates.
@@
Historical Data & Search Candidates. To evaluate the robustness of our algorithm with respect to the size of historical data, we vary the number of historical data points used in ANNS and KNN from 5000 to 25000 on RouterBench. As shown in Figure 7, the results clearly show that varying the amount of historical data used in ANNS and KNN has minimal impact on performance, with our algorithm consistently and significantly outperforming all baselines across all settings. To investigate the impact of the number of search candidates of ANNS and KNN on routing performance, we vary the candidate pool size from 3 to 10 on RouterBench. As shown in Figure 8, the results show that performance is marginally affected, with a slight improvement as the candidate pool size increases. Even with as few as 3 search candidates, it outperforms all baselines across benchmarks. Additional results on other benchmarks (see Appendix C.5) further show that our algorithm remains robust under varying historical data sizes and different numbers of ANNS/KNN candidates.
Routing Overhead & Quality of Historical Data. We evaluate the routing latency of Algorithm˜1 under varying query volumes. As shown in Appendix Table 7, our method incurs negligible decision overhead compared to existing training-free baselines. To assess robustness to data quality, we further consider a noisy setting with label perturbations and an out-of-distribution setting with distribution shift. Results in Appendix Table 8 indicate that our algorithm consistently outperforms all baselines in both settings. See Section˜C.6 for detailed results.
@@ Appendix A More Experiment Details
MMLU [hendrycks2020measuring] 14042
Hellaswag [zellers2019hellaswag] 10042
GSM8K [cobbe2021training] 7450
ARC Challenge [clark2018think] 1470
Winogrande [sakaguchi2021winogrande] 1267
MBPP [austin2021program] 427
MT-Bench [NEURIPS2023_91f18a12] 80
@@
MATH Lvl 1-5 [hendrycks2021measuring] 9884
MMLU-PRO [wang2024mmlu] 11786
GPQA [rein2024gpqa] 541
MUSR [sprague2023musr] 748
RAGBench [friel2024ragbench] 1827
Openhermes 2.5 [openhermes] 19455
@@
MMLU-PRO [wang2024mmlu] 12032
MUSR [sprague2023musr] 756
MATH Lvl 1-5 [hendrycks2021measuring] 1324
GPQA [rein2024gpqa] 1192
BBH [suzgun2022challengingbigbenchtaskschainofthought] 5761
@@
Benchmarks. We evaluate our method on three benchmarks: RouterBench (zero-shot version) [hu2024routerbench], SPROUT [somerstep2025carrot], and Open LLM Leaderboard v2 [fourrier2024open]. All three benchmarks are constructed from multiple dataset sources. Table 2 summarizes the query types in each benchmark, while Table 3 details the LLMs used and their corresponding token costs.
For RouterBench, we identify 13 different data sources in a total of 36497 samples – 6 more than those reported in the original paper – spanning a diverse range of query domains. This benchmark includes 11 different LLMs. Since it does not provide a predefined train/test split, we randomly sample 10000 queries as the test query set and treat the remaining queries as historical data. Note that we do not report per-token costs for models in RouterBench, as [hu2024routerbench] does not provide this information. Instead, the benchmark includes precomputed costs for each query across all 11 LLMs, which we directly use in our experiments.
@@
For Open LLM Leaderboard v2, we follow the same setup and data processing as [somerstep2025carrot], resulting in 21065 samples. The processed benchmark includes 5 different datasets, with each evaluating different aspects of LLM capabilities. This benchmark uses 18 different LLMs, and we randomly sample 10000 queries as the test set and use the remainder as the historical data. Since most evaluations in Open LLM Leaderboard v2 are likelihood-based, the cost is essentially determined by the length of the input. Following [somerstep2025carrot], we report the cost per input token in Table 3 and compute the total cost of each query based on input token counts.
In the main setting, the queries are embedded using bge-base-en-v1.5 [bge_embedding]. For the diversity consideration, we additionally evaluate with two different embedding models: SFR-Embedding-2_R [SFR-embedding-2] and gte-Qwen2-1.5B-instruct [li2023towards].
@@
For Model-based methods, we follow the training approach in [somerstep2025carrot] and train two separate Roberta-based models [DBLP:journals/corr/abs-1907-11692] to predict the performance and cost of each query. The performance prediction model is classification-based, aiming to select the optimal LLM with the highest expected performance, while the cost prediction model is regression-based and estimates the expected cost of an incoming query. This results in 2 different baselines:
@@
  • KNN-perf-routing [hu2024routerbench], uses KNN to select the model with the highest performance;
  • KNN-cost-routing [hu2024routerbench], uses KNN to select the model with the most available budget;
@@
We adopt HNSW [malkov2018efficient] as the main ANNS algorithm and set the number of candidate neighbors ($|R_{j}|$) to 5 for both ANNS and KNN. To assess the robustness of our algorithm under varying candidate set sizes, we further vary this number to 3, 7, and 10. However, many alternative methods, such as DiskANN [jayaram2019diskann], are interchangeable here, as listed in https://ann-benchmarks.com. For BatchSplit, we use a mini-batch size of 256 to balance LP computation cost with the low-latency requirements of the practical online routing.
@@
Optimization Implementation. We implement all optimization codes using the CVXPY [diamond2016cvxpy] package. For the one-time optimization step in our algorithm, we use the L-BFGS-B solver [zhu1997algorithm]. For BatchSplit, which involves solving a linear program for each batch, we adopt the HIGHS solver [huangfu2018parallelizing].
@@ Appendix B Theoretical Proofs
Definition 1 ($\epsilon$-Net [devanur2009adwords]).
@@ B.1 Discussion on Our Theoretical Assumptions
Random arrival order. We assume that the queries arrive in a random order, i.e., they can be picked adversarially, but their order is randomly permuted. Note that this is a weaker requirement than assuming the queries are sampled i.i.d. from an unknown distribution. This is because if the queries are sampled i.i.d., then they also satisfy the random order requirement, and thus our assumption is weaker. Furthermore, the random order model is also a popular assumption in many other online algorithms [mehta2007adwords, devanur2009adwords], since it typically allows one to go beyond worst-case hardness and obtain meaningful theoretical guarantees. In practice, we perform robustness studies by changing the query order and observe that our algorithm still outperforms baselines; see Section˜4.1.
@@
Lastly, we remark that similar reasonable lower-bound assumptions have been made in prior work on the online matching problem [goel2008online, devanur2009adwords, mehta2007adwords], where it is commonly assumed that bids for the items are small compared to the total budget.
“Estimating query features via similarity” is natural. The core intuition behind most previous literature on predictive LLM routing [hu2024routerbench, ong2025routellmlearningroutellms, wang2025mixllm, somerstep2025carrot, stripelis2024tensoropera], which trains a model-based predictor on auxiliary datasets to estimate performance and cost of future queries, aligns with Assumption 1. Specifically, these approaches assume that similar queries share similar routing-relevant features. If no such relationship exists, then historical data would be uninformative, and it would be impossible to leverage historical data for effective future routing. In this case, no meaningful inference can be made without directly querying the LLMs, and routing decisions degenerate to random selection.
One may argue that useful routing information can still be obtained by querying only a subset of LLMs. This idea is echoed in works that adopt cascading strategies, where LLMs are queried sequentially based on their observed capabilities on historical data [aggarwal2024automix, chen2023frugalgpt]. However, these approaches also fundamentally rely on a related assumption: LLMs that perform well on historical data or benchmarks are likely to perform well on future queries. Furthermore, these methods introduce significant latency and computational overhead, as they may consume substantial token budgets before reaching the optimal LLM. This makes them impractical for high-volume online routing scenarios with tight token budgets and low-latency requirements.
@@ B.3 Proof of Lemma 3
Inspired by the techniques used in [devanur2009adwords, mehta2007adwords, goel2008online], we first prove that, under the given setting, the partial performance score can be accurately estimated with high probability. Specifically, for each $i\in[M]$ and $\gamma\in\Phi$, we evaluate the following probability
@@ C.4 Total Budget
C.5 Historical Data Size & ANNS/KNN Candidates
Historical Data Size. To evaluate the robustness of our algorithm with respect to the size of historical data, we vary the number of historical data points used in ANNS and KNN from 5000 to 25000. For RouterBench, whose maximum available historical data is 26497, we randomly sample the subset of data points from 5000 to 25000. For RouterBench, which contains up to 26497 historical records, we randomly sample subsets within this range. For SPROUT, with a maximum of 30968 records, we similarly sample subsets from 5000 to 25000. For Open LLM Leaderboard v2, which has 11065 historical records, we sample subsets ranging from 5000 to 10000. As shown in Figure 13, the results clearly show that varying the amount of historical data used in ANNS and KNN has minimal impact on performance, with our algorithm consistently and significantly outperforming all baselines across all settings. This demonstrates the robustness and efficiency of our algorithm: even with as few as 5000 historical records, it remains effective for routing.
@@ C.5 Historical Data Size & ANNS/KNN Candidates
C.6 Routing Overhead & Quality of Historical Data
@@ C.7 Detailed Ablation Study
Impact of embedding models. The choice of embedding model is also important for estimating the performance score and cost of incoming queries. To evaluate the adaptivity of our algorithm to different embedding models used in ANNS, we conduct experiments with two larger embedding models: SFR-Embedding-2_R [SFR-embedding-2] and gte-Qwen2-1.5B-instruct [li2023towards]. As shown in Figure 15, our algorithm consistently outperforms all baselines across all benchmarks, demonstrating its robustness to the choice of embedding model.
@@ C.8 Routing Analysis by Query Types
To gain deeper insights into how our algorithm routes different types of queries, we analyze its routing behavior across the three benchmarks. We categorize query types based on their data sources, as each source is designed to evaluate different LLM capabilities. As shown in Figure 17, 18, and 19, our algorithm consistently achieves superior or competitive routing performance in most query types across benchmarks. These results demonstrate the generalizability and effectiveness of our algorithm in handling diverse query types, which indicates that its performance is not dependent on a narrow subset of tasks.
@@ Appendix D Query-Traffic Considerations in LLM Routing
In the practical LLM-serving systems, an additional important routing factor is query traffic patterns. However, the primary goal of this work is to develop a cost-efficient online routing algorithm, a problem that has gained increasing attention in recent literature [ong2025routellmlearningroutellms, hu2024routerbench, somerstep2025carrot], especially given the high computational cost of LLMs. To isolate this core algorithmic problem, we adopt a standard assumption widely used in the online algorithms literature [devanur2009adwords, mehta2007adwords, goel2008online]: incoming queries arrive in a randomly permuted order. This assumption, discussed in detail in Appendix B.1, allows us to abstract away complex traffic dynamics and focus on the fundamental cost-efficiency routing problem.